# Geometry Aware Operator Transformer As An Efficient And Accurate Neural Surrogate For PDEs On Arbitrary Domains

### v2

@@ Abstract
The very challenging task of learning solution operators of PDEs on arbitrary domains accurately and efficiently is of vital importance to engineering and industrial simulations. Despite the existence of many operator learning algorithms to approximate such PDEs, we find that accurate models are not necessarily computationally efficient and vice versa. We address this issue by proposing a geometry aware operator transformer (GAOT) for learning PDEs on arbitrary domains. GAOT combines novel multiscale attentional graph neural operator encoders and decoders, together with geometry embeddings and (vision) transformer processors to accurately map information about the domain and the inputs into a robust approximation of the PDE solution. Multiple innovations in the implementation of GAOT also ensure computational efficiency and scalability. We demonstrate this significant gain in both accuracy and efficiency of GAOT over several baselines on a large number of learning tasks from a diverse set of PDEs, including achieving state of the art performance on a large scale three-dimensional industrial CFD dataset. Code is available at https://github.com/camlab-ethz/GAOT.
@@ 1 Introduction
Partial Differential Equations (PDEs) are widely used to mathematically model very diverse natural and engineering systems [14]. Currently, numerical algorithms, such as the finite element and finite difference methods, are the preferred framework for simulating PDEs [44]. However, these methods can be computationally very expensive, particularly for the so-called many-query problems such as uncertainty quantification (UQ), control, and inverse problems. Here, the solver must be called repeatedly, leading to prohibitive costs and providing the impetus for the design of fast and efficient surrogates for PDE solvers [36].
To this end, ML/AI based algorithms are increasingly being explored as neural PDE surrogates. In particular, neural operators [22, 5], including those proposed in [27, 28, 33, 45, 20], which learn the PDE solution operator from data, are widely used [2]. As many of these neural operators are restricted to PDEs on Cartesian (regular) grids, they cannot be directly applied to most engineering and industrial systems, which are set on domains with complex geometries, imposing a pressing need for neural operators for learning PDEs on arbitrary domains (point clouds).
In this context, a variety of options have recently been proposed, including domain masking for FNO and CNO [45], replacing FFT in FNO with direct spectral evaluations [31], augmenting FNO with learned diffeomorphisms [26] and mapping arbitrary point cloud data between the input domain and latent regular grids with learned encoders/decoders, while processing on the latent grid with FNO [29] or transformers [1]. Alternatives such as end-to-end message-passing based graph neural networks [41, 15, 47, 46, 6, 13, 39] or end-to-end transformer based models [50, 18] have also been proposed.
@@
A thorough comparison of the existing models, not just in terms of accuracy but also computational efficiency and scalability, is necessary to evaluate whether these models are yet capable enough to act as surrogates for highly successful finite element methods for engineering simulations [44]. As one of the contributions in this paper, we performed such a comparison (see Sec.1 for details) to find evidence for an accuracy-efficiency trade-off, i.e., accurate and robust models, such as the message passing based RIGNO of [39] are not necessarily computationally efficient nor scalable in terms of training throughput and inference latency. On the other hand, more efficient models such as GINO [29] are not accurate enough (see the accompanying Radar Chart in Fig. 1). Given this observation, our main goal in this paper is to propose an algorithm to learn PDE solution operators on arbitrary domains that is accurate, computationally efficient, and can be seamlessly scaled to real-world industrial problems.
To this end, we propose Geometry Aware Operator Transformer (GAOT, pronounced goat) as a neural surrogate for PDEs on arbitrary domains. While being based on the well-established encode-process-decode paradigm [41], GAOT includes several novel features that are designed to ensure both computational efficiency and accuracy, such as
@@
  • A transformer processor that utilizes patching (as in ViTs [10]) for computational efficiency.
@@
  • A set of implementation strategies to ensure that the computational realization of GAOT is efficient and highly scalable.
@@
  • Extensively testing GAOT on 24 challenging benchmarks for both time-independent and time-dependent PDEs of various types, ranging from regular grids to random point clouds to highly unstructured adapted grids, and comparing it with widely-used baselines to show that GAOT is both highly accurate as well as computationally efficient and scalable, see Fig. 1.
  • The efficiency and scalability of GAOT is further showcased by it achieving state of the art (SOTA) performance on the large scale three-dimensional industrial benchmark of DrivAerNet++ dataset for automobile aerodynamics [12].
@@
The overall architecture of GAOT is depicted in Fig. 2. For simplicity of exposition, we start with the time-independent case, where given inputs $a(x_{j})$ on the point cloud $D_{\Delta}=\{x_{j}\}\subset D$, for $1\leq j\leq J$, GAOT provides an approximation to solution $u$ of the PDE (1) at any query point $x\in D$. At a high level, GAOT follows the encode-process-decode paradigm of [41]. In the first step, an encoder transforms the input on the underlying point cloud $D_{\Delta}$ to a latent point cloud $\mathcal{D}\subset\mathbb{R}^{d}$. The resulting spatial tokens are then processed by a processor module to learn useful representations and its output is remapped to the original domain $D$ via the decoder, which allows evaluation at any query point $x\in D$.
@@
As depicted in SM Fig. B.1, the latent domain $\mathcal{D}$ (to which the encoder maps) can be chosen in three different ways, namely i) a regular (structured) grid stencil, consisting of equispaced points on a Cartesian domain (see also Fig. 2) ii) randomly downsampling the underlying point cloud $D_{\Delta}$ or iii) a projected low-dimensional representation, where a high-dimensional domain is projected to a lower dimension (for instance using tri-plane embeddings in 3-D [8]) and a regular grid is used in the lower-dimensional domain. GAOT is a general framework where any of these latent point cloud choices can be employed for $\mathcal{D}$.
@@
Given input values $a(x_{j})$ on the underlying point cloud $D_{\Delta}$, the encoder aims to transform it into latent features $w_{e}(y)$ at any point $y\in\mathcal{D}$ on the latent point cloud. Using a graph-neural operator (GNO) encoder as in GINO [29] would lead to,
@@
The only geometric information in the afore-described encoder is provided by the coordinates of the underlying points. This alone does not convey the rich geometric information about the domain that can affect the solution of the underlying PDE (2). Hence, we need to embed further geometric information into the model. Deviating from the literature where geometric information is provided either by appending them as node and edge features on the underlying graphs [15] or by encoding a signed distance function [29], we propose to use novel geometry embeddings to encode this information. To this end and as described in SM Sec. B.3, we can rely on local statistical embeddings for each point $y\in\mathcal{D}$ as all the neighboring points $x_{k}$ in $D_{\Delta}$ with $|y-x_{k}|\leq r_{m}$ have already been computed in the AGNO encoder. From these points, we can readily compute statistical descriptors such as i) number of neighbors $x_{k}\in D_{\Delta}$, in the ball $B_{r_{m}}(y)$, ii) the average distance $D_{{\rm avg}}=\frac{1}{n^{m}_{y}}\sum_{k=1}^{n_{y}^{m}}|y-x_{k}|$, iii) the variance of this distance $D_{{\rm var}}$, with respect to the average $D_{{\rm avg}}$, iv) the centroid offset vector $\Delta_{y}=\frac{1}{n^{m}_{y}}\sum_{k=1}^{n_{x}^{y}}(x_{k}-y)$ and v) a few principal component (PCA) features of the covariance matrix of $y-x_{k}$ to calculate the local shape anisotropy. These statistical descriptors, for each scale $m$ and each point $y\in\mathcal{D}$ are then concatenated into a single vector $z_{y}$, normalized across components to yield zero mean and unit variance and fed into an MLP to provide the embedding $g^{m}(y)$. Alternatively, geometry embedding using emphPointNet models [42] can also be considered.
@@
The encoder provides a set of geometry aware tokens $w_{e}(y_{\ell})$, for all points $y_{\ell}\in\mathcal{D}$, with $1\leq\ell\leq L$, in the latent point cloud. These tokens are further transformed by a processor. As shown in Fig. 2, we choose a suitable transformer based processor. While postponing details on the processor architecture to SM Sec. B.4, we summarize our choices here. If the latent points $\{y_{\ell}\}$ lie on a regular grid (either through a structured stencil or a projected low-dimensional one), we use a patch-based vision transformer or ViT ([10] and [20, 38] for PDE operator learning) for computational efficiency. The equispaced latent points are combined into patches and the tokens in each patch are flattened into a single token embedding which serves as the input for a multi-head attention block, followed by a feed forward block. RMS normalization is applied to the tokens before processing. Either sinusoidal absolute position embeddings or rotary relative position embeddings are used to encode token positions. If the latent points $y_{\ell}$ are randomly downsampled from the original point cloud, there is no obvious way to patch them together. Hence, a standard transformer [49], but with RMS normalization, can be used. Additionally, we employ multiple skip connections across transformer blocks (see Fig. 2). The transformer processor transforms the tokens $w_{e}(y_{\ell})$ into processed tokens, that we denote by $w_{p}(y_{\ell})$, for all $1\leq\ell\leq L$.
@@
To learn the solution operator $\mathcal{S}_{t}$ of the time-dependent PDE, we observe that the $\mathcal{S}_{t}$ can be used to update the solution forward in time, given the solution at any time point $u(t)$ by applying $u(t+\tau)=\mathcal{S}_{\tau}(u(t))$. Thus, for any time $t$, given the augmented input $a(t)=(c,u(t))$, with $c$ being the coefficient in the PDE (2), we need GAOT to output $u(t+\tau)$, for any $\tau\geq 0$. To this end, we retain the architecture of GAOT, as described for the time-independent case above, and simply add the current time $t$ and the lead-time $\tau$ as further inputs to the model. More precisely, the time-dependent version of GAOT is of the form $\widehat{\mathcal{S}}_{\theta}(x,t,\tau,a(t))$, where $a(t)$ takes values at points sampled in $D$. Following [39], the map $\hat{S}_{\theta}$ can be used to update an approximate solution of PDE (2) in time by following a very general time-stepping strategy:
@@
Here, choosing the parameters $(\gamma,\delta)$ appropriately leads to different strategies for time stepping: $\gamma=0,\delta=1$ directly approximates the output of the solution operator at time $t+\tau$; $\gamma=1,\delta=1$ yields the residual of the solution at the later time, with respect to the solution at current time; $\gamma=1,\delta=\tau$ is equivalent to approximating the time-derivative of the solution. GAOT provides the flexibility to use any of these time-stepping strategies. We also use the all2all training strategy [20] to leverage trajectory data for time-dependent PDEs.
Efficient implementation.
As our goal is to ensure accuracy and computational efficiency, we have designed GAOT with ability for large-scale computations in mind. We started with the realization that the heaviest burden of the computation should fall on the processor. The encoder and decoder are often responsible for memory overheads as these modules entail sparse computations on graphs with far more edges than nodes, making the computations largely edge-based and leading to high (and inefficient) memory usage. Moreover, in many PDE learning tasks on arbitrary geometries, the underlying domain (and the resulting graph) varies significantly between data samples, making load balancing very difficult.
To address these computational challenges, we resorted to i) moving the graph construction outside the model evaluation by either storing the graph, representing the input point cloud, in memory for small graphs or on disk for large graphs and loading them during training with efficient data loaders ii) sequentially processing each input in a given batch for the encoder and decoder, while still batch processing in the transformer processor, allowing us to reduce memory usage while retaining efficiency and iii) if needed for very large-scale datasets, we use an edge-dropping strategy to further increase the memory usage of the encoder and decoder. These innovations are essential to ensure batch training and underpin the efficiency of GAOT, even when input geometries vary significantly.
@@
Time-Dependent GAOT RIGNO-18 RIGNO-12 GeoFNO FNO DSE GINO
NS-Gauss 2.91 2.29 3.80 41.1 38.4 13.1
NS-PwC 1.50 1.58 2.03 26.0 56.7 5.85
NS-SL 1.21 1.28 1.91 13.7 22.6 4.48
NS-SVS 0.46 0.56 0.73 9.75 26.0 1.19
CE-Gauss 6.40 6.90 7.44 42.1 30.8 25.1
CE-RP 5.97 3.98 4.92 18.4 27.7 12.3
Wave-Layer 5.78 6.77 9.01 11.1 28.3 19.2
Wave-C-Sines 4.65 5.35 6.25 13.1 5.52 5.82
Datasets and Baselines. We start by testing GAOT of a challenging suite of 15 datasets for PDEs with input/output data on arbitrary point clouds in two space dimensions, see SM Secs D and G. for a detailed description of the datasets and for visualizations, respectively. For time-independent PDEs, in addition to the elasticity benchmark of [26], we consider two Poisson equation datasets: Poisson-Gauss, defined on random points in a square domain, and Poisson-C-Sines, a new dataset we propose, containing rich multiscale solutions on a circular domain. In addition, we propose 4 new datasets comprising compressible flows past objects, both airfoils as well as bluff bodies. These datasets have significant variation in domain geometry and flow conditions (Mach numbers ranging from subsonic to supersonic, varying angles of attack etc.) and are tailor-made for testing neural PDE surrogates on arbitrary domains. For time-dependent PDEs, we test on the challenging datasets considered recently in [39], composed of 8 operators corresponding to the compressible Euler (2), incompressible Navier-Stokes (4), and acoustic wave equations (2). These time-dependent operators include complex multiscale solutions with shocks and other sharp traveling waves which can interact, reflect and diffract making them hard to learn. We test GAOT on these datasets and compare them with several widely used neural operators for PDEs on arbitrary domains including those based on message passing (RIGNO [39]), Fourier Layers (GINO [29], GeoFNO [26], FNO DSE[31]) and Transformers (Transolver [50], UPT [1] and GNOT [18]).
@@
It is worth reiterating that the computational efficiency of an ML model is a significant marker of overall performance. We test efficiency in terms of two critical quantities, the training throughput and the inference latency. For a given input and model size, training throughput measures the number of samples that a model can process during training (forward pass, backward pass and gradient update) per unit time (second) on a given compute system (GPU or CPU). The higher the training throughput, the faster the model can be trained. On the other hand, the inference latency is the amount of time it takes for a model to infer a single input.
We present the throughput and latency for GAOT and three selected baselines (RIGNO-18 for Graph-based, GINO for FNO-based and Transolver for Transformer-based models) in Table 2 for learning tasks (such as the bluff-body dataset for compressible flow) where the domain geometry varies throughout the dataset. These experiments are conducted on one NVIDIA-4090 GPU with float32 precision. We see from this table that GAOT has the highest training throughput and the fastest inference latency, being almost $50\%$ and $15\%$, respectively better than the second-most efficient model (GINO).
Table 2: Comparison of model size (Params.), throughtput (samples/s), and latency (ms) across GAOT and representative baselines.
@@
How the training throughput of a model scales with increasing input and model size, is absolutely crucial for evaluating whether it can be used to process large-scale datasets (input scalability) or whether it can serve as a backbone of foundation models (model scalability) which require large model sizes [20]. To evaluate the scalability of different models, we plot how the training throughput changes as input size and model size (Fig. 3 (a, b)) are increased to find that GAOT scales much more favorably than the baselines with respect to both input and model size. In fact, models like Transolver and GNOT scale very poorly, making it impossible for us to train them for the large-scale time-dependent datasets with all2all training, which requires handling large volumes of data for large input sizes. Hence, we omit them in the accuracy results for time-dependent datasets in Tab. 1. The results for both accuracy and efficiency across a range of metrics for GAOT, RIGNO, GINO and Transolver are summarized in SM Tab. E.1 and visualized in the Radar chart Fig. 1 demonstrating that GAOT ensures both accuracy (robustness) and computational efficiency (scalability) at the same time, while being the best model performing model on both sets of metrics.
The DrivAerNet++ Dataset.
Given the high accuracy and excellent computational efficiency and scalability of GAOT, we showcase its abilities further on a challenging three-dimensional large-scale benchmark for industrial simulations, the DrivAerNet++ dataset of [12]. In this benchmark, the data consists of high-fidelity CFD simulations across $8$K different car shapes which span the entire range of conventional car design. The underlying task is to learn steady-state surface fields (See Fig. 4) such as the pressure and wall shear stress, given the input car shape and flow conditions. The underlying data has approximately $500$K points per shape, making the overall training extremely compute intensive. Thus, only scalable models can currently process this learning task. We test GAOT on this challenging 3D benchmark and report the RMSE and MAE test errors for the pressure and wall shear stress in Tab. 3.
Table 3: Error Metrics of MSE (×$\times 10^{-2}$) and Mean AE (×$\times 10^{-1}$) for Pressure and Wall Shear Stress on the DrivAerNet++ Dataset.

### v3

@@ Abstract
The very challenging task of learning solution operators of PDEs on arbitrary domains accurately and efficiently is of vital importance to engineering and industrial simulations. Despite the existence of many operator learning algorithms to approximate such PDEs, we find that accurate models are not necessarily computationally efficient and vice versa. We address this issue by proposing a geometry aware operator transformer (GAOT) for learning PDEs on arbitrary domains. GAOT combines novel multiscale attentional graph neural operator encoders and decoders, together with geometry embeddings and (vision) transformer processors to accurately map information about the domain and the inputs into a robust approximation of the PDE solution. Multiple innovations in the implementation of GAOT also ensure computational efficiency and scalability. We demonstrate this significant gain in both accuracy and efficiency of GAOT over several baselines on a large number of learning tasks from a diverse set of PDEs, including achieving state of the art performance on three large scale three-dimensional industrial CFD datasets. Our project page for accessing the source code is available at camlab-ethz.github.io/GAOT.
@@ 1 Introduction
Partial Differential Equations (PDEs) are widely used to mathematically model very diverse natural and engineering systems PDEbook . Currently, numerical algorithms, such as the finite element and finite difference methods, are the preferred framework for simulating PDEs NAbook . However, these methods can be computationally very expensive, particularly for the so-called many-query problems such as uncertainty quantification (UQ), control, and inverse problems. Here, the solver must be called repeatedly, leading to prohibitive costs and providing the impetus for the design of fast and efficient surrogates for PDE solvers scimlrev2 .
To this end, ML/AI based algorithms are increasingly being explored as neural PDE surrogates. In particular, neural operators kovachki2023neural ; bartolucci2023representation , including those proposed in FNO ; GNO ; DeepONet ; CNO ; herde2024poseidon , which learn the PDE solution operator from data, are widely used azizzadenesheli2024neural . As many of these neural operators are restricted to PDEs on Cartesian (regular) grids, they cannot be directly applied to most engineering and industrial systems, which are set on domains with complex geometries, imposing a pressing need for neural operators for learning PDEs on arbitrary domains (point clouds).
In this context, a variety of options have recently been proposed, including domain masking for FNO and CNO CNO , replacing FFT in FNO with direct spectral evaluations lingsch2024beyond , augmenting FNO with learned diffeomorphisms geofno and mapping arbitrary point cloud data between the input domain and latent regular grids with learned encoders/decoders, while processing on the latent grid with FNO li2023geometryinformed or transformers UPT ; zhou2025textpde . Alternatives such as end-to-end message-passing based graph neural networks MGN ; MSMGN ; sanchez2018graph ; sanchez2020learning ; brandstetter2022message ; equer2023multi ; rigno or end-to-end transformer based models wu2024transolver ; luo2025transolver ; hao2023gnot ; wang2025cvit ; serrano2024aroma have also been proposed.
@@
A thorough comparison of the existing models, not just in terms of accuracy but also computational efficiency and scalability, is necessary to evaluate whether these models are yet capable enough to act as surrogates for highly successful finite element methods for engineering simulations NAbook . As one of the contributions in this paper, we performed such a comparison (see Sec.1 for details) to find evidence for an accuracy-efficiency trade-off, i.e., accurate and robust models, such as the message passing based RIGNO of rigno are not necessarily computationally efficient nor scalable in terms of training throughput and inference latency. On the other hand, more efficient models such as GINO li2023geometryinformed are not accurate enough (see the accompanying Radar Chart in Fig. 1). Given this observation, our main goal in this paper is to propose an algorithm to learn PDE solution operators on arbitrary domains that is accurate, computationally efficient, and can be seamlessly scaled to real-world industrial simulations.
To this end, we propose Geometry Aware Operator Transformer (GAOT, pronounced goat) as a neural surrogate for PDEs on arbitrary domains. While being based on the well-established encode-process-decode paradigm MGN , GAOT includes several novel features that are designed to ensure both computational efficiency and accuracy, such as
@@
  • A transformer processor that utilizes patching (as in ViTs dosovitskiy2021an ) for computational efficiency.
@@
  • A set of novel implementation strategies to ensure that the computational realization of GAOT is efficient and highly scalable.
@@
  • Extensively testing GAOT on 28 challenging benchmarks for both time-independent and time-dependent PDEs of various types, ranging from regular grids to random point clouds to highly unstructured adapted grids, and comparing it with widely-used baselines to show that GAOT is both highly accurate as well as computationally efficient and scalable, see Fig. 1.
  • The efficiency and scalability of GAOT is further showcased by it achieving state of the art (SOTA) performance on the large scale three-dimensional industrial benchmark of DrivAerNet++ dataset for automobile aerodynamics dvnet++ . We also test GAOT and demonstrate its superior performance to the GINO baseline on two further 3D industrial scale datasets, i.e., DrivaerML dataset for automobile aerodynamics and a NASA-CRM dataset for aerospace applications.
@@
The overall architecture of GAOT is depicted in Fig. 2. For simplicity of exposition, we start with the time-independent case, where given inputs $a(x_{j})$ on the point cloud $D_{\Delta}=\{x_{j}\}\subset D$, for $1\leq j\leq J$, GAOT provides an approximation to solution $u$ of the PDE (1) at any query point $x\in D$. At a high level, GAOT follows the encode-process-decode paradigm of MGN . In the first step, an encoder transforms the input on the underlying point cloud $D_{\Delta}$ to a latent point cloud $\mathcal{D}\subset\mathbb{R}^{d}$. The resulting spatial tokens are then processed by a processor module to learn useful representations and its output is remapped to the original domain $D$ via the decoder, which allows evaluation at any query point $x\in D$.
@@
As depicted in SM Fig. B.1, the latent domain $\mathcal{D}$ (to which the encoder maps) can be chosen in three different ways, namely i) a regular (structured) grid stencil, consisting of equispaced points on a Cartesian domain (see also Fig. 2) ii) randomly downsampling the underlying point cloud $D_{\Delta}$ or iii) a projected low-dimensional representation, where a high-dimensional domain is projected to a lower dimension (for instance using tri-plane embeddings in 3-D tripnet ) and a regular grid is used in the lower-dimensional domain. GAOT is a general framework where any of these latent point cloud choices can be employed for $\mathcal{D}$.
@@
Given input values $a(x_{j})$ on the underlying point cloud $D_{\Delta}$, the encoder aims to transform it into latent features $w_{e}(y)$ at any point $y\in\mathcal{D}$ on the latent point cloud. Using a graph-neural operator (GNO) encoder as in GINO li2023geometryinformed would lead to,
@@
The only geometric information in the afore-described encoder is provided by the coordinates of the underlying points. This alone does not convey the rich geometric information about the domain that can affect the solution of the underlying PDE (2). Hence, we need to embed further geometric information into the model. Deviating from the literature where geometric information is provided either by appending them as node and edge features on the underlying graphs MSMGN or by encoding a signed distance function li2023geometryinformed , we propose to use novel geometry embeddings to encode this information. To this end and as described in SM Sec. B.3, we can rely on local statistical embeddings for each point $y\in\mathcal{D}$ as all the neighboring points $x_{k}$ in $D_{\Delta}$ with $|y-x_{k}|\leq r_{m}$ have already been computed in the AGNO encoder. From these points, we can readily compute statistical descriptors such as i) number of neighbors $x_{k}\in D_{\Delta}$, in the ball $B_{r_{m}}(y)$, ii) the average distance $D_{{\rm avg}}=\frac{1}{n^{m}_{y}}\sum_{k=1}^{n_{y}^{m}}|y-x_{k}|$, iii) the variance of this distance $D_{{\rm var}}$, with respect to the average $D_{{\rm avg}}$, iv) the centroid offset vector $\Delta_{y}=\frac{1}{n^{m}_{y}}\sum_{k=1}^{n_{x}^{y}}(x_{k}-y)$ and v) a few principal component (PCA) features of the covariance matrix of $y-x_{k}$ to calculate the local shape anisotropy. These statistical descriptors, for each scale $m$ and each point $y\in\mathcal{D}$ are then concatenated into a single vector $z_{y}$, normalized across components to yield zero mean and unit variance and fed into an MLP to provide the embedding $g^{m}(y)$. Alternatively, geometry embedding using PointNet models Pnet can also be considered.
@@
The encoder provides a set of geometry aware tokens $w_{e}(y_{\ell})$, for all points $y_{\ell}\in\mathcal{D}$, with $1\leq\ell\leq L$, in the latent point cloud. These tokens are further transformed by a processor. As shown in Fig. 2, we choose a suitable transformer based processor. While postponing details on the processor architecture to SM Sec. B.4, we summarize our choices here. If the latent points $\{y_{\ell}\}$ lie on a regular grid (either through a structured stencil or a projected low-dimensional one), we use a patch-based vision transformer or ViT (dosovitskiy2021an and herde2024poseidon ; genCFD for PDE operator learning) for computational efficiency. The equispaced latent points are combined into patches and the tokens in each patch are flattened into a single token embedding which serves as the input for a multi-head attention block, followed by a feed forward block. RMS normalization is applied to the tokens before processing. Either sinusoidal absolute position embeddings or rotary relative position embeddings are used to encode token positions. If the latent points $y_{\ell}$ are randomly downsampled from the original point cloud, there is no obvious way to patch them together. Hence, a standard transformer vaswani2017attention , but with RMS normalization, can be used. Additionally, we employ multiple skip connections across transformer blocks (see Fig. 2). The transformer processor transforms the tokens $w_{e}(y_{\ell})$ into processed tokens, that we denote by $w_{p}(y_{\ell})$, for all $1\leq\ell\leq L$.
@@
To learn the solution operator $\mathcal{S}_{t}$ of the time-dependent PDE, we observe that the $\mathcal{S}_{t}$ can be used to update the solution forward in time, given the solution at any time point $u(t)$ by applying $u(t+\tau)=\mathcal{S}_{\tau}(u(t))$. Thus, for any time $t$, given the augmented input $a(t)=(c,u(t))$, with $c$ being the coefficient in the PDE (2), we need GAOT to output $u(t+\tau)$, for any $\tau\geq 0$. To this end, we retain the architecture of GAOT, as described for the time-independent case above, and simply add the current time $t$ and the lead-time $\tau$ as further inputs to the model. More precisely, the time-dependent version of GAOT is of the form $\widehat{\mathcal{S}}_{\theta}(x,t,\tau,a(t))$, where $a(t)$ takes values at points sampled in $D$. Following rigno , the map $\hat{S}_{\theta}$ can be used to update an approximate solution of PDE (2) in time by following a very general time-stepping strategy:
@@
Here, choosing the parameters $(\gamma,\delta)$ appropriately leads to different strategies for time stepping: $\gamma=0,\delta=1$ directly approximates the output of the solution operator at time $t+\tau$; $\gamma=1,\delta=1$ yields the residual of the solution at the later time, with respect to the solution at current time; $\gamma=1,\delta=\tau$ is equivalent to approximating the time-derivative of the solution. GAOT provides the flexibility to use any of these time-stepping strategies. We also use the all2all training strategy herde2024poseidon to leverage trajectory data for time-dependent PDEs.
Efficient implementation. As our goal is to ensure accuracy and computational efficiency, we have designed GAOT with ability for large-scale computations in mind. We started with the realization that the heaviest burden of the computation should fall on the processor. The encoder and decoder are often responsible for memory overheads as these modules entail sparse computations on graphs with far more edges than nodes, making the computations largely edge-based and leading to high (and inefficient) memory usage. Moreover, in many PDE learning tasks on arbitrary geometries, the underlying domain (and the resulting graph) varies significantly between data samples, making load balancing very difficult.
To address these computational challenges, we resorted to i) moving the graph construction outside the model evaluation by either storing the graph, representing the input point cloud, in memory for small graphs or on disk for large graphs and loading them during training with efficient data loaders ii) sequentially processing each input in a given batch for the encoder and decoder, while still batch processing in the transformer processor, allowing us to reduce memory usage while retaining efficiency and iii) if needed for very large-scale datasets, we use an edge-dropping strategy to further increase the memory usage of the encoder and decoder. These innovations are essential to ensure batch training and underpin the efficiency of GAOT, even when input geometries vary significantly. A more detailed discussion on these novel implementation tricks is provided in SM E.2.
@@
Time-Dependent GAOT RIGNO-18 GeoFNO FNO DSE UPT GINO
NS-Gauss 2.91 2.29 41.1 38.4 92.5 13.1
NS-PwC 1.50 1.58 26.0 56.7 100 5.85
NS-SL 1.21 1.28 13.7 22.6 51.5 4.48
NS-SVS 0.46 0.56 9.75 26.0 4.2 1.19
CE-Gauss 6.40 6.90 42.1 30.8 64.2 25.1
CE-RP 5.97 3.98 18.4 27.7 26.8 12.3
Wave-Layer 5.78 6.77 11.1 28.3 19.6 19.2
Wave-C-Sines 4.65 5.35 13.1 5.52 12.7 5.82
Datasets and Baselines. We start by testing GAOT of a challenging suite of 15 datasets for PDEs with input/output data on arbitrary point clouds in two space dimensions, see SM Secs D and G for a detailed description of the datasets and for visualizations, respectively. For time-independent PDEs, in addition to the elasticity benchmark of geofno , we consider two Poisson equation datasets: Poisson-Gauss, defined on random points in a square domain, and Poisson-C-Sines, a new dataset we propose, containing rich multiscale solutions on a circular domain. In addition, we propose 4 new datasets comprising compressible flows past objects, both airfoils as well as bluff bodies. These datasets have significant variation in domain geometry and flow conditions (Mach numbers ranging from subsonic to supersonic, varying angles of attack etc.) and are tailor-made for testing neural PDE surrogates on arbitrary domains in two space dimensions. For time-dependent PDEs, we test on the challenging datasets considered recently in rigno , composed of 8 operators corresponding to the compressible Euler (2), incompressible Navier-Stokes (4), and acoustic wave equations (2). These time-dependent operators include complex multiscale solutions with shocks and other sharp traveling waves which can interact, reflect and diffract making them hard to learn. We test GAOT on these datasets and compare them with several widely used neural operators for PDEs on arbitrary domains including those based on message passing (RIGNO rigno ), Fourier Layers (GINO li2023geometryinformed , GeoFNO geofno , FNO DSElingsch2024beyond ) and Transformers (Transolver wu2024transolver , UPT UPT and GNOT hao2023gnot ).
@@
Table 2: Comparison of model size (Params.), throughput (samples/s), and latency (ms) across GAOT and representative baselines.
@@
It is worth reiterating that the computational efficiency of an ML model is a significant marker of overall performance. We test efficiency in terms of two critical quantities, the training throughput and the inference latency. For a given input and model size, training throughput measures the number of samples that a model can process during training (forward pass, backward pass and gradient update) per unit time (in seconds) on a given compute system (GPU or CPU). The higher the training throughput, the faster the model can be trained. On the other hand, the inference latency is the amount of time it takes for a model to infer a single input. We present the throughput and latency for GAOT and three selected baselines (RIGNO-18 for Graph-based, GINO for FNO-based and Transolver for Transformer-based models) in Table 2 for learning tasks (such as the bluff-body dataset for compressible flow) where the domain geometry varies throughout the dataset. These experiments are conducted on one NVIDIA GeForce RTX 4090 GPU with float32 precision. We see from this table that GAOT has the highest training throughput and the fastest inference latency, being almost $50\%$ and $15\%$, respectively better than the second-most efficient model (GINO).
How the training throughput of a model scales with increasing input and model size, is absolutely crucial for evaluating whether it can be used to process large-scale datasets (input scalability) or whether it can serve as a backbone of foundation models (model scalability) which require large model sizes herde2024poseidon . To evaluate the scalability of different models, we plot how the training throughput changes as input size and model size (Fig. 3 (a, b)) are increased to find that GAOT scales much more favorably than the baselines with respect to both input and model size. In fact, models like Transolver and GNOT scale very poorly, making it impossible for us to train them for the large-scale time-dependent datasets with all2all training, which requires handling large volumes of data for large input sizes. Hence, we omit them in the accuracy results for time-dependent datasets in Tab. 1. The results for both accuracy and efficiency across a range of metrics for GAOT, RIGNO, GINO and Transolver are summarized in SM Tab. E.4 and visualized in the Radar chart Fig. 1. This demonstrates that GAOT ensures both accuracy (robustness) and computational efficiency (scalability) at the same time, while being the best model performing model on both sets of metrics.
Industrial scale 3D datasets.
Table 3: Error metrics of MSE ($\times 10^{-2}$) and Mean AE ($\times 10^{-1}$) for Pressure and Wall Shear Stress on the DrivAerNet++ dataset.